CAN SOME ONE HELP ME WITH THIS PROBLEM.
SOLVE ALGEBRAICALLY AND CHECK YOUR POTENTIAL SOLUTIONS
SQUAREROOT X+20-X=0

Answer by stanbon(75887)

SOLVE ALGEBRAICALLY AND CHECK YOUR POTENTIAL SOLUTIONS
SQUAREROOT X+20-X=0
---------------
Rewrite as:
x-sqrt(x)-20 = 0
Factor to get:
(sqrtx-5)(sqrtx+4)=0
sqrtx = 5 or sqrtx = -4
Only potential solution is sqrt(x) = 5
--------------
Checking that solution in x-sqrt(x)-20 =0
25 -5-20 = 0
20=20
==========
Cheers,
Stan H.: